OpenStudy (anonymous):

13 + (–12) – (–5) = ?

OpenStudy (anonymous):

try it 1st

OpenStudy (anonymous):

when you subtract a negative you add a postitive

OpenStudy (jiteshmeghwal9):

open the brackets 13-12+5 now use PEMDAS :)

OpenStudy (jiteshmeghwal9):

Remember:- (+)*(-)=- (-)*(+)=- (+)*(+)=+ (-)*(-)=+

OpenStudy (anonymous):

-6 ?

OpenStudy (anonymous):

try again, would it be negative?

OpenStudy (anonymous):

i think so . -4 ?

OpenStudy (anonymous):

13 -12= ?

OpenStudy (anonymous):

no, 13-12+5=?